A ferry transports tourists between three islands. It sails from the first island to the second island, 4.76 km away, in a direction 37.08 north of east. It then sails from the second island to the third island in a direction 69.08 west of north. Finally it returns to the first island, sailing in a direction 28.08 east of south. Calculate the distance between (a) the second and third islands and (b) the first and third islands.


1
Expert's answer
2020-09-29T09:39:41-0400

You may simplify the solution if draw a figure representing all the transitions:



Calculate the distance between the islands. Just break the vectors into x- and y- components:


"A_x=A\text{ cos}\alpha=3.798\text{ km},\\
A_y=A\text{ sin}\alpha=2.870 \text{ km}.\\
B_x=B\text{ sin}\beta=0.9341B,\\
B_y=B\text{ cos}\beta=0.3571B.\\
C_x=C\text{ sin}\gamma=0.4707C,\\
C_y=C\text{ cos}\gamma=0.8823C.\\"

Also, from the diagram, we see that the algebraic sum of all these vectors is zero:


"A_x-B_x+C_x=0,\\
A_y+B_y-C_y=0."

Now substitute all components with their right parts above:


"x|\space3.798-0.9341B+0.4707C=0,\\
y|\space2.870+0.3571B-0.8823C=0."

The solution of this system (and distances between is


"B=7.167\text{ km},\\
C=6.154\text{ km}."